A survey of several 9 to 11 year olds recorded the following amounts spent on a trip to the mall:

$24.73,$24.85,$22.86,$23.81,$21.31,$17.40

Construct the 80% confidence interval for the average amount spent by 9 to 11 year olds on a trip to the mall. Assume the population is approximately normal.

Step 1 of 4:

Calculate the sample mean for the given sample data. Round your answer to two decimal places.

Step 2 of 4:

Calculate the sample standard deviation for the given sample data. Round your answer to two decimal places.

Critical value?

:

Construct the 80%confidence interval. Round your answer to two decimal places.
Answer #1
Sr. X x - l2 (C- )
1 24.73 2.2367 5.0027
2 24.85 2.3567 5.5539
3 22.86 0.3667 0.1344
4 23.81 1.3167 1.7336
5 21.31 -1.1833 1.4003
6 17.4 -5.0933 25.9420
Total 134.96 0 39.7669

n = 6

Sample Mean = \frac{\sum x_{i}}{n} = 134.96 / 6 = 22.493 = 22.49

Sample SD = I - A = 39.7669 V5 = 2.8202 = 2.82

Try avoiding to round off the intermediate calculations to get more accuracy

Since we don't have the known population standard deviation we are going to use t-dist

(1 - \alpha) confidence interval for population mean

(ī Fta/2,1-17

Where critical value ta/2,n-1 = to.20/2,5 ((1 - \alpha) = 80% therefore \alpha = 0.20)

= t_{0.1,5} (using t-dist tables for df = 5 and prob = 0.1 or excel function 'tinv(0.20,5)' )

= 1.4759

Substituting the values in the interval

2.8202 (22.493371.4759 V6

(20.79.24.19)